# FastFLUX: Pruning FLUX with Block-wise Replacement and Sandwich Training

### v2

@@ 1 Introduction
Recent advances in diffusion-based generative models zhang2024attentioncalibrationdisentangledtexttoimage; esser2024scaling; podell2024sdxl; sauer2023adversarialdiffusiondistillation; chen2023pixartalphafasttrainingdiffusion; zhang2023addingconditionalcontroltexttoimage have substantially improved the fidelity and controllability of text-to-image synthesis. Transformer-based architectures such as DiT peebles2022dit and FLUX flux2024 further enhance generation quality by stacking deep hierarchical blocks that increase model expressiveness and scalability. However, these designs also introduce significant computational overhead, including large memory consumption and high inference latency, which limits their suitability for real-time generation and deployment on resource-constrained hardware.
@@
Although prior studies have investigated model compression through pruning castells2024ld; fang2023structural, knowledge distillation chen2025sana; meng2023distillation; salimans2022progressive, and quantization li2024svdquant; so2023temporal; he2023efficientdm, these approaches primarily focus on parameter-level optimization, largely overlooking the hierarchical structural organization of deep generative transformers. Specifically, hierarchical redundancy in diffusion-based generative models remains largely underexplored, despite being a significant source of inefficiency in modern deep transformer architectures. Furthermore, many existing techniques Shirkavand2025; fang2023structural require full-model fine-tuning after pruning or incur substantial performance degradation, which limits their practical applicability.
@@
In this work, we revisit the problem of hierarchical redundancy in large diffusion transformers and propose FastFLUX, a general and efficient framework for architecture-level compression. FastFLUX operates by selectively replacing the computationally heavy residual branches of FLUX blocks with lightweight linear alternatives, while preserving generation quality through a localized fine-tuning strategy. Specifically, we introduce Block-wise Replacement with Linear Layers (BRLL), which targets the high-cost residual pathways in ResBlocks. In FLUX, each ResBlock contains a heavy residual branch and a lightweight shortcut connection; the residual branch dominates both FLOPs and latency during inference. BRLL substitutes this branch with a simple linear layer while preserving the shortcut path, ensuring stable information flow. These linear layers can be initialized via least-squares estimation, providing a training-free initialization with controllable accuracy–efficiency trade-offs. To mitigate performance degradation after replacement, we propose a Sandwich Training (ST) strategy. Instead of full-model fine-tuning, ST performs localized adaptation by applying LoRA hu2021LoRAlowrankadaptationlarge modules to the neighboring unpruned layers of each replaced block, forming a small “sandwich” around the linear module. This local supervision enables the replaced branch to adapt effectively within its context, achieving stable and high-quality generation with minimal training overhead. Using these approaches, we prune 20% of the FLUX model while retaining high visual fidelity and significantly improving inference efficiency.
@@ 2 Related Work
Diffusion-based Generative Models. Diffusion models have achieved remarkable progress across diverse domains cao2024survey, establishing foundations through discrete frameworks like DDPM ho2020ddpm and latent space formulations such as Latent Diffusion rombach2022ldm. Transformer-based architectures like DiT peebles2022dit demonstrate strong scalability, with recent models such as FLUX flux2024 further increasing depth to enhance generation quality. However, deeper hierarchical stacks introduce substantial computational and memory overhead, motivating the need for efficient compression strategies.
Model Quantization and Pruning. To reduce computational costs, diffusion models employ quantization and pruning techniques shen2025efficient. Post-training quantization methods like PTQ4DM shang2023post, Q-Diffusion li2023q, and PTQD he2023ptqd address timestep-specific calibration challenges, while quantization-aware training approaches such as TDQ so2023temporal and QALoRA he2023efficientdm learn optimized representations. Complementary pruning methods remove redundant weights via Taylor expansion fang2023structural, latent space guidance castells2024ld, or dynamic programming kim2024layermerge, achieving substantial size reduction with minimal quality loss.
Transformer Compression. Transformer compression tackles computational challenges through multiple strategies tang2024survey. Knowledge distillation transfers capabilities from large teachers to compact students sanh2019distilbert; touvron2021training, while quantization reduces precision via methods like FQ-ViT lin2021fq and PTQ-ViT liu2021post. Pruning techniques eliminate redundancy through head removal yu2023x or context selection anagnostidis2023dynamic. Architectural innovations target attention mechanisms ainslie2023gqa; hatamizadeh2023fastervit, feed-forward networks du2022glam; touvron2023llama, or propose alternative designs like Mamba gu2024mamba; zhu2024vision. While these methods address parameter and structural redundancy, explicit compression of depth-wise redundancy in diffusion transformers remains underexplored.
@@ 3.2 Sandwich Training
Block Importance Estimation. We observe that different blocks contribute unequally to the model’s generative performance (see supplementary materials for more results). Based on this observation, we adopt an importance-aware pruning strategy that prioritizes the removal of less critical blocks. To quantitatively assess the importance of each block, we combine two commonly used metrics in text-to-image evaluation: FID heusel2018ganstrainedtimescaleupdate and CLIP Score radford2021learningtransferablevisualmodels. Specifically, for each ResBlock, we temporarily remove it from the model and generate a set of images. We then compute the FID and CLIP scores for each generated set and normalize both scores independently across all blocks. Since a lower FID indicates better quality, we transform it by taking $1-FID_{norm}$ so that higher values are consistently better across both metrics. The importance score $s_{i}$ for the i-th block is defined as:
@@
where $\alpha$ and $\beta$ are hyperparameters that balance the contribution of each metric. In our experiments, we set $\alpha\ =\beta\ =0.5$. A lower score $s_{i}$ implies that removing the block has a smaller impact on the overall generation quality, indicating that the block is less important and should be pruned earlier in the compression process. We therefore sort all blocks in ascending order of their importance scores and perform block replacement accordingly. Although Stable Flow StableFlow also employs layer-wise importance estimation, it relies solely on similarity-based metrics. In contrast, we incorporate both similarity and image quality assessments, enabling a more comprehensive evaluation of each block’s contribution to generation performance.
@@
DALL·E miniramesh2021zero 26.10 25.56 25.56 26.12 25.83 -
SD v1.4rombach2022highresolutionimagesynthesislatent 27.26 26.61 26.66 27.27 26.95 -
SD v2.0rombach2022highresolutionimagesynthesislatent 27.48 26.89 26.86 27.46 27.17 -
SD v3.0esser2024scaling 28.58 27.87 28.15 27.72 28.08 -
@@
DALL·E miniramesh2021zero 0.73 0.11 0.12 0.37 0.02 0.01 0.23 -
SD v1.4rombach2022highresolutionimagesynthesislatent 0.98 0.36 0.35 0.73 0.01 0.07 0.42 -
SD v2.0rombach2022highresolutionimagesynthesislatent 0.98 0.50 0.48 0.86 0.06 0.15 0.51 -
SD v3.0esser2024scaling 0.98 0.75 0.51 0.84 0.22 0.52 0.63 -
@@ 4.1 Implementation Details
Our base model is FLUX.1-dev, and all experiments are conducted using the fluxdev-controlnet-16k dataset fluxdevcontrolnet16k, which consists of approximately 16.1k images. We sample 512 images to compute the parameters of the linear layers and another 512 images for training. Since we only need to fit the parameters of a single linear layer and a few LoRA modules at a time, only a small number of samples are required for training. We set the rank to 16 and the alpha to 8 of LoRA. To provide a comprehensive and holistic assessment of our method, we utilize a range of text-to-image evaluation benchmarks, including HPS v2 wu2023humanpreferencescorev2 and GENEVAL ghosh2023genevalobjectfocusedframeworkevaluating.All experiments are implemented in PyTorch and executed on two NVIDIA RTX 3090 GPUs using a block-swap mechanism for memory-efficient training.
@@ 4.2 Performance Comparison
Quantitative Results on Multiple Benchmarks. We evaluate the T2I generation performance of our model on two widely used benchmarks: HPS v2 (for assessing alignment with complex human preferences) and GENEVAL (for evaluating short-prompt generation). The results are shown in Table1 and Table2, respectively. We compare FastFLUX with the original FLUX.1-dev, L1-norm prunedwen2016learning; michel2019sixteen variants of FLUX, and several state-of-the-art T2I modelsramesh2021zero; rombach2022highresolutionimagesynthesislatent; rombach2022highresolutionimagesynthesislatent; esser2024scaling. Since both FastFLUX and the L1-norm method are model compression techniques, we group models by similar inference latency for fair comparison. From the results, we observe that FastFLUX, when pruned by 10 percent, not only preserves generation quality but also outperforms the original FLUX.1-dev on the HPS v2 benchmark. Specifically, it achieves an improvement of 0.26 in score and ranks first among all evaluated T2I models. In addition, FastFLUX reduces inference latency by 7.48ms compared to the unpruned baseline. Within the same latency group, the L1-norm pruned model removes 12 percent of the parameters but performs 0.38 lower than FastFLUX on the HPS v2 metric. Moreover, L1-based pruning requires global fine-tuning to adapt to the irregular parameter structure, resulting in higher training costs. In contrast, FastFLUX requires only minimal localized adaptation. These observations are consistent across other latency groups and evaluation settings. Overall, FastFLUX significantly improves inference efficiency while maintaining high-quality image generation.
@@
Generalization to Other Architectures. While FastFLUX is primarily evaluated on the FLUX model with a DiT-style backbone, we further assess its generalization capability on Stable Diffusion v3.0 esser2024scaling, a model with a different architectural design. Specifically, we apply both BRLL and ST to SD v3.0, which differs in depth, attention layout, and residual structure. As shown in Table 3, the FastSD v3.0-10% variant achieves a higher CLIP Score compared to the unpruned SD v3.0 baseline, with only a moderate drop in HPS v2. These results confirm that BRLL and ST generalize well beyond DiT-style transformers and can benefit other class of diffusion-based generative models.
@@
Figure 6: Comparison of HPS v2 scores with and without linear layers, under the same training strategy. The result shows that incorporating BRLL significantly improves performance.
Figure 7: HPSv2 scores under different numbers of layers involved in ST. The x-axis indicates the total count of LoRA-enhanced layers. Performance improves substantially from 0 to 2 LoRA layers and saturates at 4, indicating a 2-layer structure provides a trade-off between performance and efficiency.
@@ 4.3 Ablation Studies
Impact of Block-wise Replacement with Linear Layer. To evaluate the necessity of block-wise replacement, we perform a variant where selected low-importance ResBlocks are directly removed without replacement. Specifically, we remove the residual branches of these blocks while preserving the shortcut connections, and apply the same ST strategy as in our full method. As shown in Figure 7, this direct removal leads to a significant drop in performance compared to using linear replacements. The HPSv2 score decreases by 0.33, from 28.58 to 28.25, confirming that the linear replacement step is crucial for preserving model capability after pruning.
@@
Number of layers in Sandwich Block. In FastFLUX, Sandwich Training applies LoRA modules to adjacent blocks around the replaced linear layer to provide localized supervision. We investigate how the number of surrounding layers involved in ST affects model performance. Specifically, a setting of 0 layers corresponds to applying no LoRA modules. A 2-layer configuration includes LoRA on $\mathbf{B}_{i-1}$ and $\mathbf{B}_{i+1}$, while a 4-layer setting extends this to $\mathbf{B}_{i-2}$, $\mathbf{B}_{i-1}$, $\mathbf{B}_{i+1}$, and $\mathbf{B}_{i+2}$. As shown in Figure 7, moving from 0 to 2 layers yields a substantial improvement in HPS v2 score by 0.61, demonstrating the benefit of local supervision from adjacent blocks. However, increasing from 2 to 4 layers results in only a marginal gain 0.03, while introducing a 25% increase in trainable parameters. This diminishing return, coupled with increased computational cost, makes the 2-layer configuration a favorable trade-off between performance and efficiency.

### v3

@@ 1 Introduction
Recent advances in diffusion-based generative models Zhang et al. (2024); Esser et al. (2024); Podell et al. (2024); Sauer et al. (2023); Chen et al. (2023); Zhang et al. (2023) have substantially improved the fidelity and controllability of text-to-image synthesis. Transformer-based architectures such as DiT Peebles and Xie (2023) and FLUX Black Forest Labs (2024) further enhance generation quality by stacking deep hierarchical blocks that increase model expressiveness and scalability. However, these designs also introduce significant computational overhead, including large memory consumption and high inference latency, which limits their suitability for real-time generation and deployment on resource-constrained hardware.
@@
Although prior studies have investigated model compression through pruning Castells et al. (2024); Fang et al. (2023), knowledge distillation Chen et al. (2025); Meng et al. (2023); Salimans and Ho (2022), and quantization Li et al. (2024); So et al. (2023); He et al. (2023a), these approaches primarily focus on parameter-level optimization, largely overlooking the hierarchical structural organization of deep generative transformers. Specifically, hierarchical redundancy in diffusion-based generative models remains largely underexplored, despite being a significant source of inefficiency in modern deep transformer architectures. Furthermore, many existing techniques Shirkavand et al. (2025); Fang et al. (2023) require full-model fine-tuning after pruning or incur substantial performance degradation, which limits their practical applicability.
@@
In this work, we revisit the problem of hierarchical redundancy in large diffusion transformers and propose FastFLUX, a general and efficient framework for architecture-level compression. FastFLUX operates by selectively replacing the computationally heavy residual branches of FLUX blocks with lightweight linear alternatives, while preserving generation quality through a localized fine-tuning strategy. Specifically, we introduce Block-wise Replacement with Linear Layers (BRLL), which targets the high-cost residual pathways in ResBlocks. In FLUX, each ResBlock contains a heavy residual branch and a lightweight shortcut connection; the residual branch dominates both FLOPs and latency during inference. BRLL substitutes this branch with a simple linear layer while preserving the shortcut path, ensuring stable information flow. These linear layers can be initialized via least-squares estimation, providing a training-free initialization with controllable accuracy–efficiency trade-offs. To mitigate performance degradation after replacement, we propose a Sandwich Training (ST) strategy. Instead of full-model fine-tuning, ST performs localized adaptation by applying LoRA Hu et al. (2022) modules to the neighboring unpruned layers of each replaced block, forming a small “sandwich” around the linear module. This local supervision enables the replaced branch to adapt effectively within its context, achieving stable and high-quality generation with minimal training overhead. Using these approaches, we prune 20% of the FLUX model while retaining high visual fidelity and significantly improving inference efficiency.
@@ 2 Related Work
Diffusion-based Generative Models. Diffusion models have achieved remarkable progress across diverse domains Cao et al. (2024), establishing foundations through discrete frameworks like DDPM Ho et al. (2020) and latent space formulations such as Latent Diffusion Rombach et al. (2022b). Transformer-based architectures like DiT Peebles and Xie (2023) demonstrate strong scalability, with recent models such as FLUX Black Forest Labs (2024) further increasing depth to enhance generation quality. However, deeper hierarchical stacks introduce substantial computational and memory overhead, motivating the need for efficient compression strategies.
Model Quantization and Pruning. To reduce computational costs, diffusion models employ quantization and pruning techniques Shen et al. (2025). Post-training quantization methods like PTQ4DM Shang et al. (2023), Q-Diffusion Li et al. (2023), and PTQD He et al. (2023b) address timestep-specific calibration challenges, while quantization-aware training approaches such as TDQ So et al. (2023) and QALoRA He et al. (2023a) learn optimized representations. Complementary pruning methods remove redundant weights via Taylor expansion Fang et al. (2023), latent space guidance Castells et al. (2024), or dynamic programming Kim et al. (2024), achieving substantial size reduction with minimal quality loss.
Transformer Compression. Transformer compression tackles computational challenges through multiple strategies Tang et al. (2024). Knowledge distillation transfers capabilities from large teachers to compact students Sanh et al. (2019); Touvron et al. (2021), while quantization reduces precision via methods like FQ-ViT Lin et al. (2021) and PTQ-ViT Liu et al. (2021). Pruning techniques eliminate redundancy through head removal Yu and Xiang (2023) or context selection Anagnostidis et al. (2023). Architectural innovations target attention mechanisms Ainslie et al. (2023); Hatamizadeh et al. (2023), feed-forward networks Du et al. (2022); Touvron et al. (2023), or propose alternative designs like Mamba Gu and Dao (2024); Zhu et al. (2024). While these methods address parameter and structural redundancy, explicit compression of depth-wise redundancy in diffusion transformers remains underexplored.
@@ 3.2 Sandwich Training
Block Importance Estimation. We observe that different blocks contribute unequally to the model’s generative performance (see supplementary materials for more results). Based on this observation, we adopt an importance-aware pruning strategy that prioritizes the removal of less critical blocks. To quantitatively assess the importance of each block, we combine two commonly used metrics in text-to-image evaluation: FID Heusel et al. (2017) and CLIP Score Radford et al. (2021). Specifically, for each ResBlock, we temporarily remove it from the model and generate a set of images. We then compute the FID and CLIP scores for each generated set and normalize both scores independently across all blocks. Since a lower FID indicates better quality, we transform it by taking $1-FID_{norm}$ so that higher values are consistently better across both metrics. The importance score $s_{i}$ for the i-th block is defined as:
@@
where $\alpha$ and $\beta$ are hyperparameters that balance the contribution of each metric. In our experiments, we set $\alpha\ =\beta\ =0.5$. A lower score $s_{i}$ implies that removing the block has a smaller impact on the overall generation quality, indicating that the block is less important and should be pruned earlier in the compression process. We therefore sort all blocks in ascending order of their importance scores and perform block replacement accordingly. Although Stable Flow Avrahami et al. (2025) also employs layer-wise importance estimation, it relies solely on similarity-based metrics. In contrast, we incorporate both similarity and image quality assessments, enabling a more comprehensive evaluation of each block’s contribution to generation performance.
@@
DALL·E miniRamesh et al. (2021) 26.10 25.56 25.56 26.12 25.83 -
SD v1.4Rombach et al. (2022a) 27.26 26.61 26.66 27.27 26.95 -
SD v2.0Rombach et al. (2022a) 27.48 26.89 26.86 27.46 27.17 -
SD v3.0Esser et al. (2024) 28.58 27.87 28.15 27.72 28.08 -
@@
DALL·E miniRamesh et al. (2021) 0.73 0.11 0.12 0.37 0.02 0.01 0.23 -
SD v1.4Rombach et al. (2022a) 0.98 0.36 0.35 0.73 0.01 0.07 0.42 -
SD v2.0Rombach et al. (2022a) 0.98 0.50 0.48 0.86 0.06 0.15 0.51 -
SD v3.0Esser et al. (2024) 0.98 0.75 0.51 0.84 0.22 0.52 0.63 -
@@ 4.1 Implementation Details
Our base model is FLUX.1-dev, and all experiments are conducted using the fluxdev-controlnet-16k dataset Nar (2024), which consists of approximately 16.1k images. We sample 512 images to compute the parameters of the linear layers and another 512 images for training. Since we only need to fit the parameters of a single linear layer and a few LoRA modules at a time, only a small number of samples are required for training. We set the rank to 16 and the alpha to 8 of LoRA. To provide a comprehensive and holistic assessment of our method, we utilize a range of text-to-image evaluation benchmarks, including HPS v2 Wu et al. (2023) and GENEVAL Ghosh et al. (2023).All experiments are implemented in PyTorch and executed on two NVIDIA RTX 3090 GPUs using a block-swap mechanism for memory-efficient training.
@@ 4.2 Performance Comparison
Quantitative Results on Multiple Benchmarks. We evaluate the T2I generation performance of our model on two widely used benchmarks: HPS v2 (for assessing alignment with complex human preferences) and GENEVAL (for evaluating short-prompt generation). The results are shown in Table1 and Table2, respectively. We compare FastFLUX with the original FLUX.1-dev, L1-norm pruned variants of FLUX, and several state-of-the-art T2I modelsRamesh et al. (2021); Rombach et al. (2022a, a); Esser et al. (2024). Since both FastFLUX and the L1-norm method are model compression techniques, we group models by similar inference latency for fair comparison. From the results, we observe that FastFLUX, when pruned by 10 percent, not only preserves generation quality but also outperforms the original FLUX.1-dev on the HPS v2 benchmark. Specifically, it achieves an improvement of 0.26 in score and ranks first among all evaluated T2I models. In addition, FastFLUX reduces inference latency by 7.48ms compared to the unpruned baseline. Within the same latency group, the L1-norm pruned model removes 12 percent of the parameters but performs 0.38 lower than FastFLUX on the HPS v2 metric. Moreover, L1-based pruning requires global fine-tuning to adapt to the irregular parameter structure, resulting in higher training costs. In contrast, FastFLUX requires only minimal localized adaptation. These observations are consistent across other latency groups and evaluation settings. Overall, FastFLUX significantly improves inference efficiency while maintaining high-quality image generation.
@@
Generalization to Other Architectures. While FastFLUX is primarily evaluated on the FLUX model with a DiT-style backbone, we further assess its generalization capability on Stable Diffusion v3.0 Esser et al. (2024), a model with a different architectural design. Specifically, we apply both BRLL and ST to SD v3.0, which differs in depth, attention layout, and residual structure. As shown in Table 3, the FastSD v3.0-10% variant achieves a higher CLIP Score compared to the unpruned SD v3.0 baseline, with only a moderate drop in HPS v2. These results confirm that BRLL and ST generalize well beyond DiT-style transformers and can benefit other class of diffusion-based generative models.
@@
Figure 6: (a) Component ablation of BRLL on FastFLUX and FLUX. w/o: FastFLUX without BRLL. Del: FLUX with block deletion. w/ F: FLUX with deleted blocks and fixed BRLL. w/ T: FLUX with deleted blocks and trained BRLL. (b) Effects of the number of LoRA-enhanced layers in ST. The x-axis indicates the total count of LoRA-enhanced layers. Performance improves substantially from 0 to 2 LoRA layers and saturates at 4, indicating a 2-layer structure provides a trade-off between performance and efficiency.
@@ 4.3 Ablation Studies
Impact of Block-wise Replacement with Linear Layer. Figure 6(a) analyzes the effect of BRLL under different ablation settings. Removing BRLL from FastFLUX causes a performance drop, with HPSv2 decreasing from 28.58 to 28.25, indicating that BRLL is essential for preserving model capacity. In contrast, directly deleting selected blocks in FLUX results in a more severe degradation. Introducing BRLL partially alleviates this issue: fixed BRLL parameters yield a moderate performance recovery, while training BRLL further improves performance. These results demonstrate that BRLL is effective and plays a crucial role in recovering model performance under structural modification.
@@
Number of layers in Sandwich Block. In FastFLUX, Sandwich Training applies LoRA modules to adjacent blocks around the replaced linear layer to provide localized supervision. We investigate how the number of surrounding layers involved in ST affects model performance. Specifically, a setting of 0 layers corresponds to applying no LoRA modules. A 2-layer configuration includes LoRA on $\mathbf{B}_{i-1}$ and $\mathbf{B}_{i+1}$, while a 4-layer setting extends this to $\mathbf{B}_{i-2}$, $\mathbf{B}_{i-1}$, $\mathbf{B}_{i+1}$, and $\mathbf{B}_{i+2}$. As shown in Figure 6(b), moving from 0 to 2 layers yields a substantial improvement in HPS v2 score by 0.60, demonstrating the benefit of local supervision from adjacent blocks. However, increasing from 2 to 4 layers results in only a marginal gain 0.03, while introducing a 25% increase in trainable parameters. This diminishing return, coupled with increased computational cost, makes the 2-layer configuration a favorable trade-off between performance and efficiency.